### v2

# Measuring Prosody Diversity in Zero-Shot TTS: A New Metric, Benchmark, and Exploration

Abstract
Prosody diversity is essential for achieving naturalness and expressiveness in zero-shot text-to-speech (TTS). However, frequently used acoustic metrics capture only partial views of prosodic variation and correlate poorly with human perception, leaving the problem of reliably quantifying prosody diversity underexplored. To bridge this gap, we introduce ProsodyEval, a prosody diversity assessment dataset that provides Prosody Mean Opinion Score (PMOS) alongside conventional acoustic metrics. ProsodyEval comprises 1000 speech samples derived from 7 mainstream TTS systems, with 2000 human ratings. Building on this, we propose the Discretized Speech Weighted Edit Distance (DS-WED), a new objective diversity metric that quantifies prosodic variation via weighted edit distance over semantic tokens. Experiments on ProsodyEval show that DS-WED achieves substantially higher correlation with human judgments than existing acoustic metrics, while remaining highly robust in speech tokenization from HuBERT and WavLM. Leveraging DS-WED, we benchmark state-of-the-art open-source TTS systems on LibriSpeech test-clean and Seed-TTS test-en, and further explorations uncover several factors that influence prosody diversity, including generative modeling paradigms, duration control, and reinforcement learning. Moreover, we find that current large audio language models (LALMs) remain limited in capturing prosodic variations. Audio samples are available at https://prosodyeval.github.io.
Index Terms—  Zero-shot TTS, prosody diversity, prosody evaluation, prosody benchmark
1 Introduction
Prosody plays a central role in spoken communication, conveying paralinguistic information such as emotion, attitude, and intent that shapes how listeners interpret meaning. Subtle variations in pitch, intensity, and duration can fundamentally alter the interpretation of an utterance even with identical text, making prosody crucial to naturalness and expressiveness in speech synthesis. While zero-shot TTS [1, 2, 3] has advanced in intelligibility and speaker similarity, prosody diversity has received far less attention. Existing evaluation practices [4] rely largely on log $F_{0}$ root mean squared error (RMSE), which correlates weakly with human judgments, captures pitch but neglects rhythm and intensity [5], and requires costly dynamic time warping (DTW). Consequently, a clear gap remains between automatic metrics and human perception of prosody diversity.
Recent efforts have shown that both continuous [6] and discrete [7] self-supervised (SSL) speech representations effectively encode prosodic information. Concurrently, NLP automatic evaluation metrics have been adapted to discrete speech representations for reference-free [8] and reference-aware [9] speech quality assessment. In particular, SpeechTokenDistance [9] is a simple approach but remains preliminary, as it focuses solely on measuring correlation with acoustic metrics without examining what the differences represent or further analyzing correlations with human perception.
With this perspective in mind, we introduce ProsodyEval, the first human-annotated dataset for prosody diversity assessment, and DS-WED (Discretized Speech Weighted Edit Distance), a new objective diversity metric that quantifies prosodic variation via weighted edit distance over semantic token sequences derived from speech SSL models such as HuBERT [10] and WavLM [11] through $k$-kmeans clustering. ProsodyEval contains 1000 synthetic samples from 7 mainstream TTS systems, paired with 2000 human ratings of prosody diversity. Experiments show that DS-WED achieves substantially higher correlation with human judgments than frequently used acoustic metrics, including log $F_{0}$ RMSE and Mel cepstral distortion (MCD), while remaining robust across models, layers, and cluster sizes. Building on DS-WED, we establish the first systematic benchmark of prosody diversity across state-of-the-art open-source TTS systems on LibriSpeech test-clean and Seed-TTS test-en. Our further explorations reveal the impact of modeling paradigms, duration control, and reinforcement learning (RL), and demonstrate that even advanced large audio language models (LALMs) like Gemini 2.5 Pro remain unreliable for prosody evaluation.
Our contributions to the community include four aspects:
  • Dataset. We present ProsodyEval, the first human-annotated dataset for prosody diversity assessment in zero-shot TTS.
  • Metrics. We propose DS-WED, a new objective prosody diversity metric based on semantic token weighted edit distance that better correlates with human ratings than existing acoustic metrics.
  • Benchmark. We establish the first benchmark of prosody diversity across state-of-the-art open-source TTS systems, offering systematic comparisons and analyses of different generative paradigms from the perspective of prosody diversity.
  • Exploration. We investigate key factors shaping prosody diversity, showing that (1) autoregressive (AR) systems outperform flow-matching based non-autoregressive (NAR) systems but not masked generative modeling (MGM); (2) duration variation is critical, yet NAR systems lack explicit duration control, and flow-matching models with implicit alignment suffer from inherent constraints; (3) RL via direct preference optimization (DPO) trades off diversity for intelligibility; and (4) current LALMs remain unreliable for prosody understanding.
  • All related resources, including data, toolkit, benchmark, and $k$-means models, will be released to facilitate future research.
2 ProsodyEval Dataset
2.1 Data Collection
We construct the ProsodyEval dataset by aggregating synthetic speech from diverse generative paradigms, spanning AR and NAR approaches, including next-token prediction [12, 13, 3], flow matching [14, 15, 16], and MGM [17]. Concretely, ProsodyEval comprises samples synthesized by seven recent open-source TTS systems, namely XTTS-v2 [12], CosyVoice [13], CosyVoice 2 [13], E2 TTS [14], F5-TTS [15], MaskGCT [17], and ZipVoice [16]. These paradigms reflect widely used approaches in modern TTS, rendering the dataset representative for evaluating prosodic variation.
Each system generates a group of five samples per input with random seeds from 0 to 4, using prompt speech and text from LibriSpeech test-clean [18] and Seed-TTS test-en [19]. To exclude synthesis errors, we filter all groups to retain only those in which every sample is subjectively perceived as word-by-word aligned with the text. In total, ProsodyEval contains 1000 synthetic speech samples.
2.2 PMOS Collection
We recruit 20 graduate students with research experience in TTS as raters and conduct a MOS test focusing on prosodic differences. In total, 2000 high-quality ratings are collected.
Evaluation Dimension   The prosodic difference score measures the extent of variation in pitch, rhythm, and stress patterns between two audio samples generated with the same model, text, and speaker. A score of 1 indicates nearly identical prosody with imperceptible variation, while a score of 5 reflects clear and consistent prosodic differences, suggesting noticeably distinct speaking styles.
Listening Test Design   All MOS tests are conducted online under controlled conditions. After a short training session, all raters are instructed to perform the assessments in a quiet environment. In each trial, a group of five audio samples generated by the same system and speaker is presented. Raters perform pairwise comparisons across all ten possible pairs within the group in random order. Each pair is rated on a five-point Likert scale according to the perceived prosodic difference. To aid judgment, both audio playback and corresponding waveform visualization are provided. Raters are encouraged to replay samples as needed, enabling them to refine their judgments and capture both subtle and pronounced differences across pairs.
2.3 Acoustic Prosody Metrics
ProsodyEval incorporates two frequently used acoustic metrics, namely log $F_{0}$ RMSE and MCD. For each pair of samples within a group, FastDTW [20], a linear-time approximation of DTW, is first applied to align their lengths, after which log $F_{0}$ RMSE and MCD are computed. Together, these two metrics serve as complementary baseline indicators for evaluating prosodic variation.
Table 1: Correlation matrix of average Pearson coefficients ($\bar{r}$) between human ratings and objective metrics, aggregated across groups via Fisher’s $Z$ transformation. Values in brackets denote 95% confidence intervals, computed in the Fisher space and back-transformed. Darker shades of green indicate stronger correlations. All correlations are statistically significant at $p<0.001$. The strongest correlation with human ratings is highlighted in bold.
PMOS log $F_{0}$ RMSE MCD DS-WED
PMOS - 0.30[0.19, 0.40] 0.66[0.58, 0.73] 0.77[0.73, 0.81]
log $F_{0}$ RMSE 0.30[0.19, 0.40] - 0.35[0.25, 0.44] 0.36[0.26, 0.45]
MCD 0.66[0.58, 0.73] 0.35[0.25, 0.44] - 0.82[0.74, 0.87]
DS-WED 0.77[0.73, 0.81] 0.36[0.26, 0.45] 0.82[0.74, 0.87] -
Table 2: Prosody diversity benchmark of zero-shot TTS systems from diverse generative paradigms on LibriSpeech test-clean and Seed-TTS test-en, using traditional acoustic metrics log $F_{0}$ RMSE and MCD as well as our proposed DS-WED, each computed as micro-average (Avg.) and rank-based score (Borda Avg.). The best results are highlighted in bold, and second-best are underlined.
System LibriSpeech test-clean Seed-TTS test-en
log $F_{0}$ RMSE MCD DS-WED log $F_{0}$ RMSE MCD DS-WED
[…]
AR (next-token prediction)
XTTS-v2 0.31 5.14 4.18 4.70 127.84 4.89 0.28 5.73 4.12 4.16 93.15 5.50
CosyVoice 0.27 3.46 4.08 4.41 120.59 4.59 0.22 3.49 4.29 4.85 75.74 4.85
CosyVoice 2 0.30 4.56 4.07 4.47 134.34 5.38 0.24 4.45 4.12 4.35 88.04 5.78
NAR (flow matching)
E2 TTS 0.27 3.26 3.40 1.87 84.91 2.11 0.20 2.78 3.33 1.57 52.35 2.18
F5-TTS 0.26 2.98 3.48 2.19 79.59 1.50 0.21 3.23 3.49 2.28 49.00 1.51
ZipVoice 0.29 4.55 3.91 3.85 114.52 3.93 0.22 3.55 3.99 4.09 58.56 2.88
NAR (masked generative modeling)
MaskGCT 0.28 4.04 4.76 6.51 139.75 5.61 0.24 4.78 5.13 6.70 80.36 5.30
3 DS-WED Metric
Given a zero-shot TTS system, we aim to quantify prosody diversity between two generated speech samples conditioned on the same text and reference speech prompt, using distinct random seeds.
Formulation   Let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ denote two synthesized speech samples. To remove the influence of leading and trailing silences, a pretrained VAD model [21] is employed to trim the raw waveforms:
$\tilde{\mathbf{X}}_{1}=\mathbf{X}_{1}[t_{1}^{\mathrm{start}}:t_{1}^{\mathrm{end}}],\qquad\tilde{\mathbf{X}}_{2}=\mathbf{X}_{2}[t_{2}^{\mathrm{start}}:t_{2}^{\mathrm{end}}],$ (1)
where $t_{1}^{\mathrm{start}}$, $t_{2}^{\mathrm{start}}$, $t_{1}^{\mathrm{end}}$, and $t_{2}^{\mathrm{end}}$ are the onset and offset timestamps predicted by the VAD model. The resulting $\tilde{\mathbf{X}}_{1}$ and $\tilde{\mathbf{X}}_{2}$ are the corresponding silence-trimmed synthesized speech samples.
For speech tokenization, we use a self-supervised speech representation model combined with $k$-means clustering to discretize the silence-trimmed speech into token sequences $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$:
$\mathbf{c}_{1}=\mathrm{Encode}_{\mathrm{spch}}(\tilde{\mathbf{X}}_{1}),\qquad\mathbf{c}_{2}=\mathrm{Encode}_{\mathrm{spch}}(\tilde{\mathbf{X}}_{2}).$ (2)
Prosodic variation is quantified as a weighted Levenshtein distance [22] between discrete speech token sequences:
$\mathrm{\text{DS-WED}}(\mathbf{c}_{1},\mathbf{c}_{2})=\min_{\pi\in\mathcal{A}(\mathbf{c}_{1},\mathbf{c}_{2})}\sum_{(i,j,o)\in\pi}w_{o}\,c_{o}(c_{1,i},c_{2,j}),$ (3)
where $\pi$ is an edit path in the alignment set $\mathcal{A}$, $o\in\{\mathrm{sub},\mathrm{ins},\mathrm{del}\}$ is the edit operation, $c_{o}(\cdot)$ denotes the corresponding edit cost, and $w_{o}$ represents an operation-dependent weight.
Discussion   DS-WED is designed to be representation-agnostic. We instantiated it with semantic tokens [23] for following reasons:
  • Semantic tokens obtained by $k$-means clustering of SSL representations from HuBERT [10] and WavLM [11] have been demonstrated to effectively capture prosodic information [7].
  • Supervised semantic tokens from S3Tokenizer [13, 3, 24] are not considered, since the introduction of sequence-level ASR loss like CTC [25] tends to distort token duration information.
  • Acoustic tokens from EnCodec [26] are not considered, as they retain low-level signal details that are not relevant to prosody.
  • Operation weights can be tuned to perceptual sensitivity. Based on listening tests showing more sensitivity to intonation and word stress than to pause duration, we increase $w_{\mathrm{sub}}$ to 1.2 while keeping $w_{\mathrm{ins}}$ and $w_{\mathrm{del}}$ at 1. The weighted edit distance can be interpreted as the minimum perceptually prosodic modifications needed to transform one speech into another at the discrete level.
4 Experiments
4.1 Correlation Analysis with Human Judgments
Setup   We evaluate correlations between human ratings and the three objective metrics DS-WED, log $F_{0}$ RMSE, and MCD on the ProsodyEval dataset across all groups. For DS-WED, speech is discretized by applying a 50-cluster $k$-means model trained on LibriSpeech 960h to the hidden embeddings from the 8th Transformer encoder layer. Since PMOS ratings are relative within groups and not comparable across groups, we compute Pearson correlation for each group and then aggregate the group-wise correlations using Fisher’s $Z$ transformation. Statistical significance is tested with a two-sided one-sample $t$-test against zero, and 95% confidence intervals are obtained in the Fisher space and back-transformed.
Results   As shown in Table 1, PMOS exhibits the strongest correlation with DS-WED ($\bar{r}=0.77$), followed by a substantial correlation with MCD ($\bar{r}=0.66$). The correlation with log $F_{0}$ RMSE is weaker but still significant ($\bar{r}=0.30$), suggesting that pitch deviations contribute to perceived differences but capture only part of the variability. Overall, these results demonstrate that DS-WED aligns most closely with subjective human judgments, substantially outperforming widely used acoustic metrics.
4.2 Efficiency Analysis
Setup   Computational efficiency is measured on ProsodyEval by Real-Time Factor (RTF), computed as processing time divided by average speech-pair duration, on NVIDIA A100 with batch size 1.
Results   Log $F_{0}$ RMSE reaches an RTF of 0.549, and MCD reaches an RTF of 0.203. Both rely on signal-processing front-ends and DTW alignment, which are CPU-bound and difficult to accelerate on GPUs. In addition, log $F_{0}$ RMSE requires mel-cepstrum computation for DTW, adding extra overhead. In contrast, DS-WED involves only a forward pass through a pretrained speech-SSL encoder followed by $k$-means clustering and edit distance at the discrete level, achieving an RTF of 0.110. It is GPU-friendly and can be further accelerated by batching. Overall, DS-WED is scalable for large-scale evaluation and practical for speech data engineering.
4.3 Ablation Studies of DS-WED
Setup   We study the effect of the SSL backbone, the encoder layer, and the number of $k$-means clusters on the correlation between DS-WED and human ratings, using the ProsodyEval dataset.
Results   Figure 1 shows that DS-WED remains quite robust across different layers, models, and vocabulary sizes, with correlations consistently around 0.7. Middle layers 6-9 achieve stronger correlations, consistent with the richer encoded prosody information. Relative smaller cluster sizes perform best, while larger ones reduce peak correlations, making the edit-distance calculation overly sensitive and misaligned with human perceptual sensitivity to prosody. Overall, WavLM-base provides more stable correlations, while HuBERT-base exhibits larger variance. The 8th layer of HuBERT-base using 50 clusters achieves the strongest correlations. We also tried the large versions, which yield slightly higher correlations but at much greater cost, hence we use the base versions for all experiments.
Fig. 1: Ablation of average correlations between DS-WED and human ratings across models, layers, and cluster sizes on ProsodyEval.
4.4 Benchmarking Zero-shot TTS Systems
System Details   We evaluate the following open-source zero-shot TTS systems across three representative paradigms: next-token prediction, flow matching, and MGM.
Setup   Our evaluations are conducted on two widely used benchmarks: professionally read audiobooks LibriSpeech test-clean and crowdsourced read speech Seed-TTS test-en. To measure prosody diversity, we report two conventional acoustic metrics, log $F_{0}$ RMSE and MCD, together with our proposed DS-WED. Each metric is computed in two ways: (1) micro average across all samples, and (2) rank-based Borda aggregation, where systems are ranked within each group and assigned scores from seven for the best system to one for the worst, and the average score across groups is reported, thereby eliminating the influence of absolute metric values.
Results   As shown in Table 2, AR systems consistently outperform NAR flow-matching systems in prosody diversity. However, when compared with NAR MGM systems, AR systems show comparable performance. MaskGCT even surpasses all AR systems on LibriSpeech and remains competitive on Seed-TTS.
Table 3: Effect of duration perturbation (DP) on prosody diversity of NAR TTS systems on LibriSpeech test-clean and Seed-TTS test-en, using DS-WED, computed as micro-average (Avg.).
System LibriSpeech test-clean Seed-TTS test-en
$\uparrow DS-WED Avg.\uparrow$
F5-TTS 79.59 49.00
F5-TTS w/ DP 100.88+26.7% 62.95+28.5%
MaskGCT 139.75 80.36
MaskGCT w/ DP 159.10+13.8% 92.71+15.4%
Table 4: Effect of DPO on prosody diversity of zero-shot TTS systems on LibriSpeech test-clean and Seed-TTS test-en, using DS-WED, computed as micro-average (Avg.).
System LibriSpeech test-clean Seed-TTS test-en
$\uparrow DS-WED Avg.\uparrow$
CosyVoice 2 134.34 88.04
CosyVoice 2 w/ DPO 109.09-18.8% 71.64-18.6%
MaskGCT 139.75 80.36
MaskGCT w/ DPO 135.75-2.9% 77.80-3.2%
Table 5: Average Pearson ($\bar{r}$) correlations between LALM-as-Judges and prosody-related metrics, aggregated across groups via Fisher’s $Z$ transformation. Values in brackets denote 95% confidence intervals, computed in the Fisher space and back-transformed. * marks correlations statistical significance at $p<0.05$.
PMOS log $F_{0}$ RMSE MCD DS-WED
Gemini 2.5 Pro 0.27*[0.16, 0.38] 0.10[-0.08, 0.27] 0.16*[0.01, 0.30] 0.22*[0.05, 0.38]
4.5 Insights from Benchmarking and Further Exploration
Do AR zero-shot TTS systems generate more prosody diversity than NAR systems? AR systems indeed outperform flow-matching models in prosody diversity but offer no advantage over MGM.
  • AR models generate speech sequentially, providing explicit temporal modeling and natural variation in duration. In contrast, NAR flow-matching systems such as E2-TTS, F5-TTS, and ZipVoice pursue architectural simplicity by adopting implicit alignment. When regression objectives are applied to these weakly aligned and entangled representations, the models collapse toward mean predictions of highly multimodal and diverse prosodic patterns, resulting in blurry and over-smoothed outputs [28]. Average upsampling-based alignment combined with a text encoder in ZipVoice partly alleviates this issue, but the problem remains non-trivial. By comparison, MGM introduces stochasticity through iterative mask-and-prediction.
  • All NAR systems are trained on the same 100k-hour Emilia corpus, yet MaskGCT performs best in DS-WED while F5-TTS ranks lowest. Moreover, XTTS-v2, trained on less than one-third of the data, still shows superior prosody diversity, suggesting that modeling paradigm rather than data scale is the dominant factor.
To what extent does duration variation shape prosody diversity in NAR zero-shot TTS systems? We apply duration perturbation (DP) with factors 0.8, 0.9, 1.0, 1.1, and 1.2 to audio samples within each group. Prior work [17, 29] shows duration variations in this range have little effect on intelligibility. As shown in Table 3, DP consistently increases prosody diversity for two NAR systems, indicating that duration variation during inference is a key factor shaping prosody. Notably, F5-TTS exhibits a relative increase of nearly 30% with DP, yet still lags behind AR and MGM systems without DP, suggesting that prosodic monotony in flow-matching systems with implicit alignment stems from inherent architectural limitations.
Does reinforcement learning (RL) reduce prosody diversity in zero-shot TTS systems? We evaluate the prosody diversity of zero-shot TTS systems from [30], which are aligned via Direct Preference Optimization (DPO) on the INTP dataset [30] for intelligibility, using vanilla DPO for AR and extended DPO [30] for NAR MGM. As shown in Table 4, applying DPO consistently reduces prosodic diversity for AR and NAR systems, reflecting the general tendency of RL to prune variation while amplifying reward-aligned behaviors.
Can large audio language models (LALMs) serve as reliable evaluators of prosodic differences? We test Gemini 2.5 Pro [31] by prompting it to rate the relative prosodic difference between two samples within a group of five. It listens to all five samples to form a reference range of variation and then assigns a score from 1 to 5. As shown in Table 5, Gemini’s scores show a statistically significant but weak correlation with human ratings, while correlations with objective metrics fluctuate with wide confidence intervals. Combined with high prompt sensitivity, these findings suggest that Gemini 2.5 Pro is not a reliable evaluator of prosodic variation.
5 Conclusion
In this work, we introduce ProsodyEval, the first human-annotated dataset for assessing prosody diversity in zero-shot TTS, together with DS-WED, a new objective diversity metric based on weighted edit distance over semantic tokens. DS-WED outperforms frequently used acoustic metrics in correlating with human judgments and remains robust in speech discretization across models, layers, and cluster sizes, enabling more reliable evaluation of prosodic variation. Leveraging DS-WED, we systematically benchmark and analyze mainstream zero-shot TTS systems. Our findings indicate that (1) AR systems outperform NAR flow-matching models but not MGM, (2) duration variation strongly shapes prosody diversity, yet NAR systems lack duration control, and flow-matching models with implicit alignment suffer from inherent architectural limitations, (3) RL via DPO improves intelligibility while reducing prosody diversity, and (4) even advanced LALMs like Gemini 2.5 Pro remain unreliable for prosody evaluation. One limitation lies in the cross-lingual applicability of DS-WED, validated only on English. Looking ahead, ProsodyEval and DS-WED fill a gap in zero-shot TTS evaluation and open new avenues for speech data engineering.